OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9 8 − 4x 8 + 4x −8 + 4x −8 − 4x

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@mathstudent55

OpenStudy (aivantettet26):

C

OpenStudy (anonymous):

ok thsnk you

OpenStudy (aivantettet26):

oh wait

OpenStudy (aivantettet26):

yes its C

OpenStudy (mathstudent55):

We need to solvve the first equation for y: 4x - y = 8 Subtract 4x from both sides: -y = -4x + 8 Divide both sides by -1: y = 4x - 8 or y = -8 + 4x Option C

OpenStudy (anonymous):

Systems of equations with different slopes and different y-intercepts have no solutions. Always Sometimes Never -------------------------------------------------------------------------------- Question 3 (Multiple Choice Worth 4 points) [6.01] Systems of equations have one solution. Always Sometimes Never

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Never Sometimes